# Character is Destiny: Can Large Language Models Simulate Persona-Driven Decisions in Role-Playing?

### v1

@@ Abstract
Can Large Language Models (LLMs) substitute humans in making important decisions? Recent research has unveiled the potential of LLMs to role-play assigned personas, mimicking their knowledge and linguistic habits. However, imitative decision-making necessitates a more nuanced understanding of personas. In this paper, we benchmark the ability of LLMs in persona-driven decision-making. Specifically, we investigate whether LLMs can predict characters’ decisions provided with the preceding stories in high-quality novels. Leveraging character analyses written by literary experts, we construct a dataset LifeChoice comprising 1,401 character decision points from 395 books. Then, we conduct comprehensive experiments on LifeChoice, with various LLMs and methods for LLM role-playing. The results demonstrate that state-of-the-art LLMs exhibit promising capabilities in this task, yet substantial room for improvement remains. Hence, we further propose the CharMap method, which achieves a 6.01% increase in accuracy via persona-based memory retrieval. We will make our datasets and code publicly available.
Figure 1: An example of NextDecisionPrediction. LLMs generate a profile and memory from the historical data before the decision node in the storyline and then make life choices from the perspective of the character under the current scenario.
@@ 1 Introduction
If we can thoroughly understand all the causes, the future is not an unpredictable miracle but an inevitable result of the past.
– Voltaire
Decision-making is a vital part of human activities, and every important decision reflects a person’s history. The recent progress of large language models (LLMs)OpenAI (2023); Touvron et al. (2023); Team (2023) have enabled them to simulate the personas of assigned individuals, inheriting knowledge, linguistic patterns, and even personalities Park et al. (2023); Gao et al. (2024); Wang et al. (2024). However, a key challenge towards persona-playing LLMs lies in persona-driven decision-making, i.e., whether LLMs could make decisions consistent with the intended personas, as depicted by their past behaviors. This capability is indispensable for a broad range of applications, such as understanding human behavior patterns Fellows (2004), evaluating role-playing models Wang et al. (2023, 2024), and constructing personal assistant Xu et al. (2022); Salemi et al. (2024).
Previous work has studied language models in understanding characters in fictional works, e.g., guessing a character’s identities  (Sang et al., 2022) or possible actions (Mostafazadeh et al., 2016) from the context. However, their studies are limited to short and straightforward contexts. In contrast, recent persona-assigned LLMs are built on massive persona data, making their decision-making process much more complicated. Previous analysis in persona-assigned LLMs generally focuses on linguistic styles (Wang et al., 2023), knowledge memorization (Shao et al., 2023) and personality traits (Wang et al., 2024). These studies are restricted in evaluating surface-level behaviors, leaving it unclear whether persona-assigned LLMs can replicate the thought process and mindset for well-considered actions, i.e., their decision-making. Hence, the persona-driven decision-making of LLMs remains an important yet underexplored question.
In this work, we propose NextDecisionPrediction, a novel task to benchmark the ability of LLMs in simulating the decision-making process of the target personas. As shown in 1, each sample in NextDecisionPrediction is a key decision point of a character, with character data preceding the point as the input and their decision as ground truth for the output. In delicately plotted novels, the key character decisions are typically foreshadowed, leading to their destiny. Therefore, we construct LifeChoice, a multi-choice dataset based on literature analysis for fictional characters, contributed by literature experts from the Supersummary [^1] https://www.supersummary.com/. The construction of LifeChoice includes three steps: decision point detection, distractor generation, and motivation extraction. To predict the decision, LLMs need to identify and reason over the persona’ historical data related to the current scenario.
We evaluate persona-assigned LLMs on LifeChoice in various settings. Following prior research (Li et al., 2023; Wang et al., 2023), we first craft character profiles containing descriptions and memories, which compress the preceding character data. Then, we build persona-assigned LLMs based on character profiles. With extensive experiments, we find that existing persona-assigned LLMs achieve an accuracy up to 76.95% on LifeChoice, and both well-summarized character descriptions and accurate memory retrieval are important to persona-assigned LLMs. Furthermore, we propose CharMap, which adopts persona-based memory retrieval and gains a 6.01% accuracy improvement over the baselines.
@@
  • We propose NextDecisionPrediction, the first task to evaluate the decision-making capabilities of persona-assigned LLMs.
  • We construct LifeChoice, a dataset for NextDecisionPrediction. It is constructed from life choices of characters in literature work, sourced from literature analysis from experts. Besides, We propose CharMap, which adopts persona-based memory retrieval to improve LLM role-playing.
  • With extensive experiments, we find that state-of-the-art persona-assigned LLMs exhibit promising capabilities in persona-driven decision-making. Additionally, we analyze methods and future directions for further improvement.
@@ 2 Related Work
Early research on character-related studies focuses on character understanding. Brahman et al. (2021) attempts to predict a specific character through the text of the novel. Yu et al. (2022) provides dialogues from movie scripts for the model to examine and then asks it to identify the character who speaks each passage. With the enhancement of model abilities, some work attempts to make the model simulate complex role-playing. Li et al. (2023) analyzes 32 anime characters using 54k dialogues and personality traits. They use sentence embeddings for dialogue selection and evaluation. Zhou et al. (2023) uses identity, interests, and relationships, collecting AI behaviors for imitation and using character data for fine-tuning. They evaluate model consistency and linguistic style. Wang et al. (2023) creates a dataset for script characters and evaluates role-playing quality based on speaking style imitation and role-specific knowledge. These studies make a chatbot for a certain character, but they focus more on imitating the character from the perspective of dialogue, which is a shallow imitation. We aim to role-play from the perspective of behavior and decision-making. This form tests the model’s understanding of the role more.
@@
3 Datsets and Task Setups
@@ 3.1 Dataset Construction
Input:
In 1815 Monsieur Charles-François-Bienvenu Myriel was Bishop of Digne. He was then……Jean Valjean had retaken his seat at the end of the table. He had resumed his thinking. His reflections gave him a sort of frightening aspect. He was subject to one of those violent inner tearings, which was not unknown to him.
@@
In the courtroom, an innocent man was wrongfully accused of being him because he bore a resemblance to Jean Valjean. If Jean Valjean did not come forward, this innocent man would be sent to the gallows in his place. At this time, Jean Valjean had transformed his identity and become a respected town mayor, and he had also adopted a young girl named Cosette, with whom he had a new life.
@@
We build our dataset on the Supersummary website, which provides high-quality summaries and analyses of novels by literary experts. We connect with this website and utilize their full book summaries and character analyses. With these resources, we design life-choice questions for key characters using GPT-4. The dataset construction comprises the following three main steps:
Selecting Available Novels
To prevent data leakage, we filter all novels on the site using the following criteria: (1) The narrative must exclude non-fiction genres, such as biographies or documentary literature. (2) The narrative perspective must be in the first or third person. (3) The progression of narrative time should be linear, avoiding stories with complex timelines or flashbacks. For each book that passes these filters, we apply GPT-4 to parse the character analysis provided by literary experts, extracting important decision nodes and their motivations. If these motivations relate to events from earlier chapters, we identify the original chapter text through comprehensive summaries.
Constructing Multiple-Choice QA Data
We input the full novel summaries, characters’ decision nodes, and potentially influential chapter texts into GPT-4. Our goal is to generate multiple-choice questions that capture the complexity of the characters’ decision-making processes. Every option is designed to seem plausible, while only one choice—the one made in the original book—is correct.
Manual Filtering
We invite ten native English-speaking university students to filter all the data. We supply the annotators with each data sample’s corresponding summaries and human analyses, enabling them to determine whether the multiple-choice questions created by the model are challenging and reasonable. The detailed annotation rules are available in Appendix B.1.
In the end, we collect 1,401 characters from 396 books and their corresponding life choices. Table 1 shows a complete data example.
@@ 3.2 Dataset Analysis
We refer to the drama theory of Aristophanes Sommerstein (2013); Silk (2002) as the system prompt, use GPT-4 to classify the motivations for character decisions into two meta-motivations and several accompanying sub-motivations:
@@
Plot-driven behavior stems from a series of external events and the unfolding of conflicts. Characters often react passively within a larger narrative structure, with their actions led by external events. Sub-motivations of plot-driven behavior include External Conflicts, Tasks and Goals, Puzzles and Secrets, Pursuits and Escapes, Exploration and Discovery, Power and Control, and Intrigue and Betrayal.
Note that each topic will only be assigned one category of motivation. Figure 2 shows the proportion of different motivations. Detailed introductions for each sub-motivation are in Appendix A.2.
@@ 3.3 Task setups
Given an input $\mathcal{X}=(\mathcal{D},\mathcal{S},\mathcal{C},\mathcal{Q},\mathcal{A})$, comprising the preceding textual content from the original book before the decision node $\mathcal{D}$, the current scenario $\mathcal{S}$, the focused character $\mathcal{C}$, a multiple-choice question $\mathcal{Q}$ outlining a decision faced by that character, and a set of candidate answers $A=\{A_{i}\}_{i=1}^{4}$. The ultimate goal is to identify the correct choice $\mathcal{Y}$ that aligns with the character’s decision in the narrative.
This task can be formulated as $P(\mathcal{Y}|\mathcal{X})$. For evaluation, we directly use the accuracy of multiple-choice QA. As shown in Table 2, compared to other character understanding tasks, the LifeChoice requires understanding the characters through an excessively long context to make selections. Similar to Personal LLM assistants, the model needs to locate relevant information in a vast amount of sparse personal data related to the current scene. This behavior demands a more profound understanding of the characters.
@@ 4 Experiments
Profile Construction Reasoning Model Task QA Profile Score
ACC +motivation
Only description
Model DescriptionWu et al. (2021) gpt-4-1106-preview 48.43 89.24 1.84
Human Description gpt-4-1106-preview 60.02 95.47 2.23
Only memory
Model Retrieval(BM25) gpt-4-1106-preview 26.08 86.88 1.07
Model Retrieval(Embedding) gpt-4-1106-preview 38.66 89.24 1.75
Description+Memory
Model Description+Model Retrieval gpt-3.5-turbo-1106 51.62 96.39 1.96
gpt-4-1106-preview 58.92 93.46 2.05
@@
gpt-4-1106-preview 76.95 100.00 2.67
Table 3: Results of different LLMs performance on NextDecisionPrediction. ACC refers to the decision accuracy. +motivation refers to the results after adding manually written motivation by experts to the constructed profile, and Profile Score refers to the score given by GPT-4 for whether the profile contains the meaning of motivation.
Since our inputs generally exceed 100k, it is difficult for LLMs to handle them directly. Therefore, our method is divided into two steps: First, we compress the input to under 5k tokens(the maximum token count for human description) to construct a character profile that complies with most LLMs’ token restrictions. Second, we feed the constructed profile and the question into different LLMs for responses.
@@ 4.1 Character Profile Construction
For the first step, we refer to the role-playing model Li et al. (2023); Zhou et al. (2023) and propose two baseline methods. The first method summarizes all the context into a description of the corresponding character, composed of the character’s basic situation and storyline. This provides a holistic portrayal of the character during role-playing. The second method focuses on retrieving memories of the current scene, which could offer more detail. Below, we detail the methods we use:
@@
We adopt the method by Wu et al. (2021) to summarize the characters’ profiles. Considering the model’s capabilities and costs, we use GPT-3.5 to replace GPT-3 in the original paper. We divide $\mathcal{D}$ into chunks, summarize them separately, and then merge them. After several iterations, we generate the final summary. We also use the manually written character storylines in Supersummary as the human description.
@@
We use two retrieval methods to find memory from $\mathcal{D}$ relevant to the current scenario $\mathcal{S}$. The first is the BM25 algorithm Robertson et al. (2009), and the second uses embedding-based retrieval. For the embedding retrieval, we break $\mathcal{D}$ into chunks of 1k tokens and match them with vectors extracted by the embedding model for $f(\mathcal{S},\mathcal{Q})$. Both retrieval results were kept within 5k tokens. For the embedding model, we use OpenAI’s text-embedding-ada-002Neelakantan et al. (2022) model.
@@
Using description and memory alone may result in missing information thus we attempt to concatenate the results of both methods. For the merged text, we keep it within 5k tokens.
4.2 Reasoning the Answer
After compressing the original input $\mathcal{X}$ into a Character Profile, we feed it into the LLMs. For the individual methods, we respectively use GPT-4 as the final reasoning model. For the joint method, we experiment with the following LLMs: Mixtral-8x7B-MoE Jiang et al. (2024), Claude2.1 Anthropic (2023), GPT-3.5-Turbo OpenAI (2022), GPT-4-Turbo OpenAI (2023). For all these models, we adopt the official instruction formats where available.

### v2

# Character is Destiny: Can Role-Playing Language Agents Make Persona-Driven Decisions?

@@ Abstract
Can Large Language Models (LLMs) simulate humans in making important decisions? Recent research has unveiled the potential of using LLMs to develop role-playing language agents (RPLAs), mimicking mainly the knowledge and tones of various characters. However, imitative decision-making necessitates a more nuanced understanding of personas. In this paper, we benchmark the ability of LLMs in persona-driven decision-making. Specifically, we investigate whether LLMs can predict characters’ decisions provided by the preceding stories in high-quality novels. Leveraging character analyses written by literary experts, we construct a dataset LifeChoice comprising 1,462 characters’ decision points from 388 books. Then, we conduct comprehensive experiments on LifeChoice, with various LLMs and RPLA methodologies. The results demonstrate that state-of-the-art LLMs exhibit promising capabilities in this task, yet substantial room for improvement remains. Hence, we further propose the CharMap method, which adopts persona-based memory retrieval and significantly advances RPLAs on this task, achieving 5.03% increase in accuracy. Resources are available at https://github.com/airaer1998/LifeChoice.
Figure 1: An example of LifeChoice. Given a character, a decision point and the preceding context, RPLAs are expected to reproduce the original decision. Typically, RPLAs are constructed by parsing the context into the character’s description and memory.
@@ 1 Introduction
The fault, dear Brutus, is not in our stars, but in ourselves, that we are underlings.
– Julius Caesar. Act 1, Scene 2.
With the recent advancements in large language models (LLMs) OpenAI (2023); Touvron et al. (2023), Role-Playing Language Agents (RPLAs) have emerged as a flourishing field of AI applications and research Chen et al. (2024). RPLAs are LLM-based AI systems that simulate assigned personas, reproducing their tones, knowledge, personalities and even decisions Park et al. (2023); Gao et al. (2024); Wang et al. (2024); Xie et al. (2024). They emulate various characters across extensive applications, including fictional characters in chatbots and video games Wang et al. (2023, 2024), as well as digital clones (Gao et al., 2023) or personalized assistants Xu et al. (2022); Salemi et al. (2024) for real-world individuals.
Can RPLAs reliably make decisions that align with their personas, as humans do? This question is vital for the practical usage of RPLAs, yet remains underexplored. Previous studies primarily investigate RPLAs’ character fidelity in terms of their tones (Wang et al., 2023) and knowledge (Shao et al., 2023), which could be readily replicated by existing RPLAs via style imitation and knowledge retrieval. However, these features are relatively superficial compared with the underlying thinking and mindset of characters. Recent efforts (Wang et al., 2024) study the personality fidelity of RPLAs, but they fail to capture the nuances and dynamics of characters’ mindsets. Hence, it remains an understudied question whether RPLAs could simulate persona-driven decisions, which challenges their comprehensive understanding of the personas and reasoning about unobserved behaviors.
In this paper, we systematically study the capability of RPLAs to simulate persona-driven decisions, based on characters from high-quality novels. In high-quality novels, characters’ life choices are carefully plotted and aligned with their personas. Hence, we introduce the LifeChoice dataset, which evaluate whether RPLAs can faithfully reproduce the characters’ life choices in the narratives. Specifically, LifeChoice comprises 1,462 character decisions from 388 novels, leveraging expert-written character analyses. Each sample is presented as a multiple-choice question with the preceding context before the decision point. As depicted in Figure 1, RPLAs are expected to identify and reason over relevant knowledge about the characters to simulate their decisions. The construction of LifeChoice primarily involves three steps: decision point selection, multiple-choice question construction, and manual examination.
Compared with previous methods for RPLA evaluation, our task and dataset benefit from higher-quality data and are more challenging. First, our questions and decisions are well-designed and closely aligned with the personas, since they are sourced from well-crafted narratives. Hence, our data establish solid ground truth for simulating characters’ persona-driven decisions. Second, our task is more challenging as it requires RPLAs to comprehensively understand and reason based on the personas, including their knowledge, experiences, and personalities. Specifically, LifeChoice poses the following challenges: 1) Long-context understanding, where RPLAs need to identify sparse relevant motivations from massive character contexts. 2) Temporal intelligence, where RPLAs should intelligently adapt to the dynamic evolution of characters and environments. 3) Intricate motives, where RPLAs are required to reason through complex and entangled backgrounds and motives to arrive at the decisions.
We conduct extensive experiments to evaluate RPLAs on LifeChoice. Our experiments cover various LLMs and different RPLA frameworks, including memory-enhanced agents, long-context LLMs, and our proposed method CharMap towards better simulation of persona-driven decisions. The results demonstrate that existing RPLAs have shown a promising accuracy of up to 62.92% on LifeChoice. Furthermore, CharMap significantly enhances RPLAs on this task, achieving an accuracy of 67.95%, which exceeds previous baselines by 5.03%. However, compared to the human performance of 92.01%, there is still significant room for improvement. Meanwhile, we observe that both well-summarized character descriptions and accurate memory retrieval are crucial for RPLAs.
@@
  • We propose to explore RPLAs’ ability in simulating persona-driven decisions, which is crucial for future RPLA applications and challenges existing RPLAs.
  • We delicately craft LifeChoice, the first benchmark for persona-driven decisions of RPLAs, based on characters’ life choices from high-quality novels. Besides, we propose CharMap, which adopts persona-based memory retrieval for better decision-making of RPLAs.
  • Based on LifeChoice, we conduct extensive experiments. The results demonstrate the promising performance of RPLAs in decision simulation. Then, we analyze and compare methodologies for RPLA development, and show the effectiveness of CharMap.
@@ 2 Related Work
Early research on character-related studies focuses on character understanding. Brahman et al. (2021) attempts to predict a specific character through the text of the novel. Yu et al. (2022) provides dialogues from movie scripts for the model to examine and then asks it to identify the character who speaks each passage. With the enhancement of model abilities, some work attempts to make the model simulate complex role-playing. Li et al. (2023) analyzes 32 anime characters using 54k dialogues and personality traits. They use sentence embeddings for dialogue selection and evaluation. Zhou et al. (2023) uses identity, interests, and relationships, collecting AI behaviors for imitation and using character data for fine-tuning. They evaluate model consistency and linguistic style. Wang et al. (2023) creates a dataset for script characters and evaluates role-playing quality based on speaking style imitation and role-specific knowledge. These studies make a chatbot for a certain character, but they focus more on imitating the character from the perspective of dialogue, which is a shallow imitation. Our study aims to evaluate RPLAs from the perspective of behavior and decision-making, which further challenges the model’s understanding of the roles.
@@
3 Dataset and Task Setups
@@ 3.1 Dataset Construction
Context:
In 1815 Monsieur Charles-François-Bienvenu Myriel was Bishop of Digne. He was then……Jean Valjean reflections gave him a sort of frightening aspect. He was subject to one of those violent inner tearings, which was not unknown to him.
@@
In the courtroom, an innocent man was wrongfully accused, because he bore a resemblance to Jean Valjean. If Jean Valjean did not come forward, this innocent man would be sent to the gallows in his place. At this time, Jean Valjean had transformed his identity and become a respected town mayor, and he had also adopted a young girl named Cosette, with whom he had a new life.
@@
We construct a comprehensive dataset called LifeChoice. As shown in Table 1, the sample for each decision point includes the preceding context $p$ from the original book, the current scenario $s$, a question $q$ outlining a decision faced by that character $c$, a list of options $a=\{a_{i}\}_{i=1}^{4}$, the correct answer $y$, and the motivation $m$ explaining the character’s choice. Our data is sourced from the website Supersummary [^1] https://www.supersummary.com/, which provides three pieces of content written by literary experts: key character descriptions, chapter summaries, and book analyses. We contact the website and obtain authorization to use the data for academic research. The dataset construction comprises the following three main steps:
Selecting Decision Points
To prevent data leakage, we first filter novels on the site using the following criteria: (1) The narrative must exclude non-fiction genres like biographies or documentary literature. (2) The narrative perspective must be in the first or third person. (3) The progression of narrative time should be linear, avoiding stories with complex timelines or flashbacks. (4) Exclude overly popular books, as measured by a high number of reviews on literary review websites. For each book that passes these filters, we provide GPT-4 with content written by literary experts, requesting it to output each key character’s life choice decision points and their corresponding gold motivations. Additionally, we also ask GPT-4 to output the chapter numbers related to the decision based on the extracted motivations. As shown in the example in Figure 1, the literary expert’s analysis of the book suggests that Michael Corleone’s motivation for choosing to assassinate the enemy includes both avenging his father and witnessing the collusion between the police and the enemy, which exposes him to the darker side of the government. We then identify two corresponding chapters in the original book based on these motivations, providing more refined data for constructing multiple-choice questions.
Constructing Multiple-Choice Questions
We input the content written by literary experts and the corresponding chapters identified based on motivation into GPT-4. Our goal is to generate multiple-choice questions that capture the complexity of the characters’ decision-making processes. The correct option reflects the decision made by the characters in the original books, whereas the distractors are designed to be plausible for an arbitrary person. As shown in the example in Figure 1, Michael Corleone can ask for help from the government because he was once a Navy officer who trusted the government. However, in the preceding text, Michael witnesses the dark side of the government, so he ultimately chooses to stab the police.
Manual Examination
We invite ten native English-speaking university students to filter the data and pay them according to local minimum wage standards. We supply the annotators with content written by literary experts and the multiple-choice questions, asking them to assess whether the model-created questions are challenging and reasonable. They are also tasked with filtering out data they deem low quality.
Ultimately, we collect 1,462 characters from 388 books and their corresponding life choices. Table 1 shows a complete data example. The specific prompts and more detailed data construction process can be found in Appendix A.
@@ 3.2 Dataset Analysis
We refer to the drama theory of Aristophanes Sommerstein (2013); Silk (2002) as the system prompt and use GPT-4 to classify the motivations for character decisions into two meta-motivations and several accompanying sub-motivations:
@@
Plot-driven behavior stems from a series of external events and conflicts unfolding. Characters often react passively within a larger narrative structure, with their actions led by external events. Sub-motivations of plot-driven behavior include External Conflicts, Tasks and Goals, Puzzles and Secrets, Pursuits and Escapes, Exploration and Discovery, Power and Control, and Intrigue and Betrayal.
Note that each topic is assigned one category of motivation. Figure 2 shows the proportion of different motivations. Detailed introductions for each sub-motivation are in Appendix B.2.
@@ 3.3 Task Setups
This task can be formulated as $P(y|x)$. Given the input $x=(p,s,c,q,a)$, the RPLA needs to identify the correct choice $y$ that aligns with the character’s decision in the narrative. For evaluation, we directly use the accuracy of multiple-choice question answering. As shown in Table 2, compared to other character understanding tasks, LifeChoice requires understanding the character through a more extended context to make decisions. RPLAs must locate relevant information related to the current scene in vast personal data. This behavior demands a more profound understanding of the characters.
@@ 4 Experiments
Profile Construction Role-Playing Model ACC +motivation
Description Construction
Hierarchical merging LLaMA-3 42.10 83.09
GPT-3.5 39.85 80.00
GPT-4 45.43 85.24
@@
GPT-3.5 52.04 86.33
GPT-4 55.17 90.23
Memory Retrieval
BM25 GPT-4 26.08 75.88
Embedding GPT-4 35.66 78.24
Description & Memory
Direct concatenation LLaMA-3 57.02 92.04
Mixtral 58.56 91.75
Claude-3.5 62.85 96.45
Gemini-1.5-pro 57.16 91.38
GPT-3.5 55.62 90.39
GPT-4 62.92 95.46
[…]
Mixtral 65.02 92.05
Claude-3.5 67.13 96.90
Gemini-1.5-pro 63.94 91.39
GPT-3.5 61.62 90.95
GPT-4 67.95 96.87
Table 3: Results of different LLMs on LifeChoice. ACC refers to the decision accuracy. +motivation refers to the results after providing the motivations behind character decisions, which are extracted from expert analyses by GPT-4.
Because our inputs generally exceed 100k, it is difficult for LLMs to handle them directly. Therefore, our approach is divided into two steps: 1) Character Profile Construction, which includes the character’s description and memories; 2) Reasoning for Decisions, where different LLMs use the constructed profile to answer the questions.
@@ 4.1 Character Profile Construction
As shown in Figure 1, the character profile consists of two parts. The first part is the character’s description, including their personality, experiences, hobbies, etc. The second part is the character’s memories, specific segments from the preceding text. Below, I will detail the methods for constructing these two parts:
@@
We adopt two automatic methods to construct character descriptions: (1) Hierarchical merging Wu et al. (2021): Books are divided into chunks that fit within the LLM context window. The LLM summarizes each chunk, then merges and summarizes adjacent summarized chunks iteratively to produce the final description. (2) Incremental updating Chang et al. (2023): Books are divided into chunks and summarized sequentially, and the description is updated and refined incrementally by concatenating summarized chunks. The summarization model for both automated methods is GPT-3.5. Additionally, using the (3) expert-written descriptions from Supersummary, we employ GPT-4 to identify the positions of the decision points and truncate the text, providing only the data before these points. All descriptions are kept within 5k tokens, the maximum for human-written descriptions.
@@
We use two memory retrieval methods: (1) BM25 Robertson et al. (2009): Scores documents based on term relevance and length, optimizing retrieval using term frequency and distribution. (2) Embedding-based retrieval: Uses dense vectors representing documents and queries to assess semantic similarity through vector distance. For the embedding model, we use OpenAI’s text-embedding-ada-002Neelakantan et al. (2022) model.
Figure 3: An overview of CharMap, a two-step scenario-specific character profile building approach.
@@
Using only Description or Memory alone may lead to information loss (Wang et al., 2024). Therefore, we also experiment by combining the results of both methods to form the character’s profile. We adopt two methods: (1) Direct concatenation: This method concatenates the results from both approaches by prompting the model to role-play the corresponding character. By default, it uses the results from Human Description and Embedding retrieval. (2) CharMap: To better utilize the information in the Description, we propose CHARacter MAPping Profile Synthesis (CharMap), constructing a more scenario-specific profile in two steps. As shown in Figure 3, first, after obtaining the description, we input it along with the question into the model, asking it to locate the plot in the Description relevant to the current scene based on the question. Second, we use these episodes as queries to retrieve related memories and then input them into the inference model and the description. This leverages the overall character storyline in the description, thereby better retrieving related memories.
4.2 Reasoning for Decisions
After compressing the original input $x$ into a character profile, we feed it into the LLMs. For methods using only description or memory, we use GPT-3.5, GPT-4, and LLaMA-3Team (2024b). For methods using both, we also include Claude-3Anthropic (2024), GeminiTeam (2024a), and Mixtral Jiang et al. (2024). For all these models, we adopt the official instruction formats where available  [^2] The versions in this paper are gpt-3.5-turbo-1106, gpt-4-1106-preview, Llama-3-70B-Instruct, Claude -3.5-Sonnet, Gemini-1.5-pro and Mixtral-8x7B-v0.1 respectively..